Grace read one quarter of a book on Monday. Tuesday she read one quarter
of the remaining pages. Wednesday she read one quarter of the pages she had
not yet read. Thursday she read the final 81 pages. How many pages are in the
book?
(A) 192 (B) 243 (C) 256 (D) 324 (E) 446

Answer by mananth(16949)

Let the number pages in the book be x.
day............. read...........balance................. Read next day
..
mon............. x/4 .......... x-x/4 = 3x/4............. 3x/16
..
tues............ 3x/16..........3x/4-3x/16= 9x/16........ 9x/64
..
wed.............9x/64.
..
Thur.............81
--------
x/4 +3x/16+9x/64 +81= x
x/4 +3x/16+9x/64- x= -81
LCM 64
16x+12x+4x-64x / 64 = -81
-27x= -81*64
x= 192
The book has 192 pages.